2x + 3y = 1
5x + 3y = 16
x = ?
y = ?

Answer by tutorcecilia(2144)

-5(2x + 3y = 1)[Use the elimination method to solve for x and y]
+2(5x + 3y = 16)
---------------
.
.
-10x-15y=-5
+10x+6y=32
-------------
+0x-9y=27
.
.
-9y=27 [Solve for y]
-9y/-9=27/-9
y=-3 [Plug (y=-3) back into the equations and solve for x]
.
2x + 3y = 1
2x + 3(-3) = 1
2x+(-9)=1
2x=9+1
2x=10
x=10/2
x=5
.
Plug the values of x and y back into either of the original equations.
5x + 3y = 16
5(5) + 3(-3) = 16
-9=-9 [Checks out]
or
2x + 3y = 1
2(5) + 3(-3) = 1
10-9=1
1=1 [Also checks out]